FIRST OBSERVATION OF CO TRIMER AND A NEW LOOK AT CO DIMER

M. REZAEI, S. SHEYBANI-DELOUI, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, AB T2N 1N4, Canada; K.H. MICHAELIAN, Natural Resources Canada, CanmetENERGY, 1 Oil Patch Drive, Suite A202, Devon, AB T9G 1A8, Canada; \underline{A.R.W. McKELLAR}, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada.

A broad-band (2135 to 2200 cm^{−1}) infrared spectrum is recorded using a tunable quantum cascade laser to probe a supersonic jet expansion of a dilute mixture of carbon monoxide in helium, giving an effective rotational temperature of about 2.5 K. Structure in the region of 2144 cm^{−1} is assigned to the previously elusive CO trimer.^{a} In spite of interference from the CO dimer and some remaining unexplained details, there is strong evidence for a planar, cyclic, C-bonded trimer structure, with C_{3h} symmetry and 4.42 Å intermolecular separation, in agreement with theoretical calculations.^{b} A modest vibrational blue-shift of +0.85 cm^{−1} is observed for the trimer, as compared to +0.71 cm^{−1} for the C-bonded form of the dimer (and −1.28 cm^{−1} for the O-bonded form). Analysis of the new dimer spectrum reveals the first known excited state (v_{CO} = 1) levels with A^{+} symmetry, and establishes that resonant vibrational splittings are small (<0.2 cm^{−1}) for both the C-bonded and O-bonded dimer isomers. The dimer spectrum extends over a surprisingly large range, with somewhat reduced intensity above 2150 cm^{−1}. A total of 28 new rotational level stacks having A^{−} symmetry are assigned for v_{CO} = 1 on the basis of combination differences, adding to the 8 stacks previously known,^{c} and extending up to 51 cm^{−1} above the v_{CO} = 1 origin. For the first 13 stacks we can establish a correspondence between v_{CO} = 0 and 1, and give labels in terms of values for K, for the low frequency intermolecular geared-bend mode, and for the intramolecular mode (in phase or out of phase C−O stretch).

aM. Rezaei et al., J. Chem. Phys. 138, 071102 (2013).

bG.W.M. Vissers et al., J. Chem. Phys. 122, 054306 (2005); T.A. Ford, Spectrochim. Acta A 64, 1151 (2006).

cM.D. Brookes and A.R.W. McKellar, J. Chem. Phys. 111, 7321 (1999); K.A. Walker et al., J. Chem. Phys. 113, 6618 (2000).